A uniform 1.2-kg rod that is 0.96 m long is suspended at rest from the ceiling by two springs, one at each end. Both springs hang straight down from the ceiling. The springs have identical lengths when they are unstretched. Their spring constants are 22 N/m and 55 N/m. Find the angle that the rod makes with the horizontal.
Answer #1

Given data,

Spring constants, k1 = 55 N/m

Spring constant k2 = 22 N/m

Length = 0.96 m

Mass , m = 1.2 kg

As the rod is in equilibrium the net torque acting on the rod is 0

On the right side of the rod,

m*g*L/2 = F*L

where, F is the restoring force

m*g / 2 = k1*x1

1.2 * 9.8 / 2 = 55 * x1

x1 = 0.106m

On the left side of the rod

m*g*L/2 = k2*x2*L

1.2 x 9.8 / 2 = 22 x x2

x2 = 0.267 m

Now, dx = x2 - x1

= 0.267 - 0.106

= 0.161 m

sin \theta = dx / L

sin \theta = 0.161 / 0.96

sin \theta = 0.1677

\theta = sin-1 (0.1677)

= 9.6540

Therefore, the angle is 9.6540